# Compactly supported $\mathbb{A}^{1}$-Euler characteristics of symmetric powers of cellular varieties

### v1

@@ 2.2. Power structures
In this section, we give a brief introduction to the power structures studied by Gusein-Zade, Luengo and Melle-Hernández in [GZLMH06] and by Pajwani and Pál in [PP23]. Informally, a power structure on a ring $R$ is a way to make sense of the expression $f(t)^{r}$ for $r\in R$ and $f(t)\in 1+tR[[t]]$, see e.g. [PP23, Definition 2.1] for a precise definition. By [GZLMH06, Proposition 1], under some finiteness assumptions it suffices to define $(1-t)^{-r}$ for $r\in R$ satisfying some conditions which we specify now.
@@
Let $R$ be a ring. A finitely determined power structure on $R$ is a collection of functions $a_{i}:R\to R$ for $i\in\mathbb{Z}_{\geq 0}$ such that:
@@
    $a_{n}(r+s)=\sum_{i=0}^{n}a_{i}(r)a_{n-i}(s)$ for all $r,s\in R$.
For the purposes of this paper, all power structures will be finitely determined. Suppose $R$ and $S$ are rings with power structures on them given by functions $a_{i}$ and $b_{i}$ respectively, and let $f:R\to S$ be a ring homomorphism. Then we say that $f$ respects the power structures if $f(a_{i}(r))=b_{i}(f(r))$ for all $i>0$ and $r\in R$.
@@
Gusein-Zade, Luengo and Melle-Hernández [GZLMH06, Page 3] proved that there is a canonical power structure on $\operatorname{K}_{0}(\mathrm{Var}_{k})$, given on the level of quasiprojective varieties by functions $S_{n}$ such that $S_{n}([X])=[X^{(n)}]$. Their paper works over base field $\mathbb{C}$, however the construction works over a general base field of characteristic zero.
This paper is concerned with the following power structure on $\operatorname{GW}(k)$ from [PP23, Corollary 3.25].
@@
For every $n\geq 0$, define functions $a_{n}:\operatorname{GW}(k)\to\operatorname{GW}(k)$ such that for $\alpha\in k^{\times}$
@@
where $t_{\alpha}=\langle 2\rangle+\langle\alpha\rangle-\langle 1\rangle-\langle 2\alpha\rangle$. Note that $t_{\alpha}$ is $2$-torsion in $\operatorname{GW}(k)$. These functions uniquely define a power structure on $\operatorname{GW}(k)$ by [PP23, Corollary 3.25].
In fact it is shown that if there is a power structure $b_{n}$ on $\operatorname{GW}(k)$ such that $\chi^{\mathrm{mot}}(X^{(n)})=b_{n}(\chi^{\mathrm{mot}}(X))$ for $X=\operatorname{Spec}(L)$ where $L/k$ is a quadratic étale algebra, then $b_{n}=a_{n}$ for all $n$. The power structure on $\operatorname{GW}(k)$ given by these $a_{n}$ functions is therefore of interest for computing the compactly supported $\mathbb{A}^{1}$-Euler characteristic of symmetric powers of varieties, since it predicts that $\chi^{\mathrm{mot}}(X^{(n)})=a_{n}(\chi^{\mathrm{mot}}(X))$ for every variety $X/k$. It is currently an open question whether this is true for all varieties, however [PP23, Corollary 4.30] shows that this is true whenever $X$ is dimension 00. We extend this result to $\operatorname{K}_{0}$-étale linear varieties in Theorem 4.10.
@@
We have that $t_{\alpha}=0$ if and only if $[\alpha]\cup[2]=0\in H^{2}_{\mathrm{Gal}}(k,\mathbb{Z}/2\mathbb{Z})$, so in particular, $t_{1}=t_{-1}=0$. Corollary 3.28 of [PP23] therefore tells us if [2]$\cup$[2]- ∪ [ 2 ] is the zero map, then $a_{n}(\langle\alpha\rangle)=\langle\alpha^{n}\rangle$ for all $n$. In particular, the power structure defined by the $a_{n}$ functions will agree with the non factorial symmetric power structure on $\operatorname{GW}(k)$ as defined by McGarraghy in [McG05, Definition 4.1].
We will use the following results about this power structure later.
Lemma 2.10.
Let $m\in\mathbb{Z}$ and let $n$ be odd. Then $a_{n}(m\mathbb{H})$ is hyperbolic.
Proof.
Since $m\mathbb{H}=m\langle 1\rangle+m\langle-1\rangle$ and $t_{1}=t_{-1}=0$, we see $a_{n}(m\mathbb{H})=S_{n}(m\mathbb{H})$, where $S_{n}$ is the non factorial symmetric power structure on $\operatorname{GW}(k)$ as in [McG05, Definition 4.1] so when $m\geq 0$ this follows by [McG05, Corollary 4.13]. For $m<0$,
$0=a_{n}(m\mathbb{H}+(-m)\mathbb{H})=\sum_{i=0}^{n}a_{i}(m\mathbb{H})a_{n-i}((-m)\mathbb{H}),$
so the result follows by induction using that $q\mathbb{H}=\mathrm{rank}(q)\cdot\mathbb{H}$ for any $q\in\operatorname{GW}(k)$. ∎

### v2

@@ 2.2. Power structures
In this section, we give a brief introduction to the power structures studied by Gusein-Zade, Luengo and Melle-Hernández in [GZLMH06] and by the first author and Pál in [PP25]. Informally, a power structure on a ring $R$ is a way to make sense of the expression $f(t)^{r}$ for $r\in R$ and $f(t)\in 1+tR[[t]]$, see e.g. [PP25, Definition 2.1] for a precise definition. By [GZLMH06, Proposition 1], under some finiteness assumptions it suffices to define functions $b_{n}:R\to R$ for $r\in R$, which should be thought of as defining $(1-t)^{-r}:=\sum_{n\geq 0}b_{n}(r)t^{n}$, satisfying some conditions which we specify now.
@@
Let $R$ be a ring. A finitely determined power structure on $R$ is a collection of functions $b_{n}:R\to R$ for $n\in\mathbb{Z}_{\geq 0}$ such that:
@@
    $b_{n}(r+s)=\sum_{i=0}^{n}b_{i}(r)b_{n-i}(s)$ for all $r,s\in R$.
For the purposes of this paper, all power structures will be finitely determined. Suppose $R$ and $S$ are rings with power structures on them given by functions $b_{i}$ and $b_{i}^{\prime}$ respectively, and let $f:R\to S$ be a ring homomorphism. Then we say that $f$ respects the power structures if $f(b_{i}(r))=b_{i}^{\prime}(f(r))$ for all $i>0$ and $r\in R$.
@@
Gusein-Zade, Luengo and Melle-Hernández [GZLMH06, Page 3] proved that there is a canonical power structure on $\mathrm{K}_{0}(\mathrm{Var}_{k})$, given on the level of quasiprojective varieties by functions $S_{n}$ such that $S_{n}([X])=[X^{(n)}]$. Their paper works over base field $\mathbb{C}$, however the construction works over a general base field of characteristic zero. In characteristic $p\neq 2$ Bejleri–McKean proved in [BM25, Lemma 6.7] that the same is true in positive characteristic if we instead work in $\mathrm{K}^{uh}_{0}(\mathrm{Var}_{k})$
This paper is concerned with the following power structure on $\operatorname{GW}(k)$ from [PP25, Corollary 3.26].
@@
For $n\geq 0$, define functions $a_{n}:\operatorname{GW}(k)\to\operatorname{GW}(k)$ such that for $\alpha\in k^{\times}$
@@
where $t_{\alpha}=\langle 2\rangle+\langle\alpha\rangle-\langle 1\rangle-\langle 2\alpha\rangle$. Note that $t_{\alpha}$ is $2$-torsion in $\operatorname{GW}(k)$. These functions uniquely define a power structure on $\operatorname{GW}(k)$ by [PP25, Corollary 3.26].
In fact it is shown in [PP25, Corollary 3.14] that if there is a power structure $b_{n}$ on $\operatorname{GW}(k)$ such that $\chi_{c}^{\mathrm{mot}}(X^{(n)})=b_{n}(\chi_{c}^{\mathrm{mot}}(X))$ for $X=\operatorname{Spec}(L)$ where $L/k$ is a quadratic étale algebra, then $b_{n}=a_{n}$ for all $n$. In [PP25, Corollary 3.27], it is deduced that if there exists a power structure $b_{n}$ on $\operatorname{GW}(k)$ such that $\chi_{c}^{\mathrm{mot}}(X^{(n)})=b_{n}(\chi_{c}^{\mathrm{mot}}(X))$ for all $n$ and all varieties $X$, i.e. if there exists a power structure $b_{n}$ on $\operatorname{GW}(k)$ such that if we give $\mathrm{K}_{0}(\mathrm{Var}_{k})$ the power structure given by symmetric powers, then $\chi_{c}^{\mathrm{mot}}$ respects the power structures, then $b_{n}$ is necessarily equal to $a_{n}$. The power structure on $\operatorname{GW}(k)$ given by these $a_{n}$ functions is therefore of interest for computing the compactly supported $\mathbb{A}^{1}$-Euler characteristic of symmetric powers of varieties, since if $\chi_{c}^{\mathrm{mot}}$ would respect power structures (with symmetric powers as the power structure on $\mathrm{K}_{0}(\mathrm{Var}_{k})$ and $a_{n}$ on $\operatorname{GW}(k)$), this would mean that $\chi_{c}^{\mathrm{mot}}(X^{(n)})=a_{n}(\chi_{c}^{\mathrm{mot}}(X))$ for every variety $X/k$. It is currently an open question whether $\chi_{c}^{\mathrm{mot}}$ does actually respect the power structures, however [PP25, Corollary 4.30] shows that this is true whenever $X$ is dimension 0. We extend this result to $\mathrm{K}_{0}$-étale linear varieties in Theorem 4.10.
@@
We have that $t_{\alpha}=0$ if and only if $\alpha]\cup[2]=0\in H^{2}_{\mathrm{Gal}}(k,\mathbb{Z}/2\mathbb{Z})$, so in particular, $t_{1}=t_{-1}=0$. One way to see this, is to use that $t_{\alpha}=-(\langle 1\rangle-\langle 2\rangle)(\langle 1\rangle-\langle\alpha\rangle)$ is a product of Pfister forms, so under the isomorphism from the Milnor conjectures, we have that $t_{\alpha}$ is mapped to $\alpha]\cup[2]$. Alternatively, an elementary proof of this statement is provided in [PP25, Lemma 3.29]. Therefore if $\cup[2]$ is the zero map, then $a_{n}(\langle\alpha\rangle)=\langle\alpha^{n}\rangle$ for all $n$. In particular, the power structure defined by the $a_{n}$ functions will then agree with the non factorial symmetric power structure on $\operatorname{GW}(k)$ as defined by McGarraghy in [McG05, Definition 4.1].
We will use the following elementary results about this power structure later.